1) Assume that a country's production function is Y = AK 0.3 L 0.7 (and MPK = 0.3 Y/K ) The ratio of capital to output is 3, the growth rate of output is 3 percent, and the depreciation rate is 4 percent. Assume the economy is in a steady state.

a.Write down the steady state condition and calculate the saving rate for this steady state.

b.Write down the Golden Rule for this economy. Is this economy in the Golden Rule steady state?

c.What saving rate would be required to reach the Golden Rule steady state?
Answer #1

1) Steady state condition, is a state investment is equal depreciahon, where where equal to KETI KE(i+n) (+7) Keel Now (las) Kat Byt KROKEE RA Kt Pulting (tn) (147) Kt =(1-6) Kt tsit Ke Ultnt + St-1) = SYE Keinth + 8) = 14 K3 x 10.07): & To.21 s wee b) on the golden hele find sawings rule which maximizer steady state level or growth of consumptionTo reach the golden rule steady stale =) nt nts should be equal to 0.1 (MPK) k* A rt tts Putting (m+ K +6 = 0.1, Pu 3. s 0.1